Consider a competitive market served by many domestic and foreign firms. The domestic demand for these firms’ product is Qd = 800 - 1.5P. The supply function of the domestic firms is QSD = 150 + 0.5P, while that of the foreign firms is QSF = 250.

Instructions: Enter your responses for equilibrium price rounded to the nearest penny (two decimal places). Enter your responses for equilibrium quantity rounded to one decimal place.

a. Determine the equilibrium price and quantity under free trade.

Equilibrium price: $

Equilibrium quantity: ? units

b. Determine the equilibrium price and quantity when foreign firms are constrained by a 100-unit quota.

Equilibrium price: $

Equilibrium quantity: ? units
Answer #1

Answer:

Given

Total Demand Qd=800-1.5P

Domestic supply Qsd=150+0.5P

Foreign supply Qf=250

Total Supply Qs=Qsd+Qf=150+0.5P+250=400+0.5P

A)

So for equilibrium Qd=Qs

800-1.5P=400+0.5P

2P=400

Equilibrium price P=$200

equilibrium quantity Qd=800-1.5*200=500 units

B)

Given

Total Demand Qd=800-1.5P

Domestic supply Qsd=150+0.5P

New Foreign supply Qf=100

Total Supply Qs=Qsd+Qf=150+0.5P+100=250+0.5P

So for equilibrium Qd=Qs

800-1.5P=250+0.5P

2P=550

Equilibrium price P=$275

equilibrium quantity Qd=800-1.5*275=387.5 units